# Light-controlled strong coupling of optical cavity modes spaced by 200 THz

### v1

@@ I Introduction
Figure 1: Optical pump-mediated coupling of cavity modes at 780 nm and 516 nm. (a) The experimental cavity: an AR-coated and 5% MgO doped periodically-poled lithium niobate (ppLN) crystal is placed within a macroscopic Fabry-Pérot cavity which is doubly-resonant at 780 nm (difference-frequency or DF) and 516 nm (sum-frequency or SF) wavelengths. The 10 mm long crystal is positioned approximately at the center of the 82 mm long cavity (SI A) (b) When pumped at 1529 nm, the crystal couples the two cavities modes ($\hat{a}_{\text{DF}}$ and $\hat{a}_{\text{SF}}$) via sum- and difference-frequency generation processes. When the pump is near two-mode resonance, this coupling admixes the two modes yielding eigenmodes (with lowering operators $\hat{a}_{+}$ and $\hat{a}_{-}$) that are the superposition of the DF and SF modes. In the absence of loss, the minimum frequency splitting between the hybridized modes is determined by the coupling strength $g=\sqrt{\epsilon}\leavevmode\nobreak\ \nu$, where $\epsilon$ is the single-pass wavelength-conversion efficiency and $\nu$ is the cavity free-spectral range ($c/2L_{\text{cav}}$). (c) We measure the resonator transmission at DF and SF as a function of the detuning of the DF probe laser from the DF cavity resonance and observe a mode splitting in the spectrum. The frequency splitting is evidence of hybridization of the DF and SF modes. The solid lines are fits. The spectra correspond to a coupling $g=2\pi\times 5.0(2)$ MHz and cavity decay rates $\{\kappa_{\text{DF}},\kappa_{\text{SF}}\}=2\pi\times\{2.5(1),6(1)\}$ MHz, providing a cooperativity $\mathcal{C}=4g^{2}/\kappa_{\text{DF}}\kappa_{\text{SF}}=7(1)$.
@@ V Acknowledgments
This work was supported by AFOSR grant FA9550-22-1-0279, ARO grant W911NF-23-1-0053, and AFOSR MURI Grant FA9550-19-1-0399. We thank Tingran Wang and Zeyang Li for assistance with instrumentation, Ash Kumar for discussion of theory, and Zeyang Li and Adam Shaw for feedback on the manuscript.
@@ X Additional Information
Each horizontal slice on the avoided-crossing diagram in the main text figure 2 is the output of the cavity when it is probed with 780 nm light. At the two-mode resonance, we measure two peaks with equal heights. Away from resonance, we measure more 780 nm transmission corresponding to the mode that has more contribution from the uncoupled DF eigenmode. The near-Lorentzian measurement of the off-resonant peaks is truncated in transmission due to the vertical resolution of the oscilloscope.
@@
Supplemental Information Fig S3: Expected improvements (a) Calculation for the single-pass efficiency per mW of pump power for the pump and input beam sizes for the current cavity design (with 780 nm cavity waist $\sim$110 $\mu$m at DF), intermediate waist ($\sim$50 $\mu$m) and optimal waist ($\sim$14 $\mu$m). The maximum is obtained when the parameters $\zeta$ of the cavity and the pump are equal to 2.84, where $\zeta=L_{\text{crystal}}/2z_{R}$ and $z_{R}$ is the Rayleigh length of the pump and the cavity. The calculations are done for 1 mW of pump power using the expression in SI C. (b) Expected end-to-end conversion efficiency in different scenarios. Blue: fit corresponding to data in main text Fig 3. Orange: with optimized outcouplers with the current crystal and cavity/pump waists. Green: with optimized waists and outcouplers with the current crystal. Red: with optimized waists and outcouplers and 3x lower losses at both SF and DF. Purple: with 10x lower losses at both SF and DF. (c) Finesse measurements with an intracavity undoped, unpoled 3 mm long lithium niobate crystal provided a limiting cavity finesse of 9900 calculated from cavity finesse without crystal (7400) and cavity finesse with crystal (4200).
@@
Other crystals, such as potassium titanyl phosphate (KTP), exhibit lower absorption losses than lithium niobate (LN). However, due to their lower electro-optic coefficients (around half of LN for KTP), they require a longer interaction length. After accounting for the length, ppKTP should provide a factor of $\sim 14$ improvement in loss over 5%-MgO doped LN at 780 nm, but evaluates to comparable overall losses as compared to congruent LN (Loss of ppKTP at 775 nm has been measured to be $\sim$127 ppm per cm steinlechner2013Absorption , $\sim$3435 ppm per cm for 5%-MgO doped LN schwesyg2011Optical and 400 ppm per cm for congruently grown LN leidinger2015Comparative ).

### v2

@@ I Introduction
Figure 1: Optical pump-mediated coupling of cavity modes at 780 nm and 516 nm. (a) The experimental cavity: an AR-coated and 5% MgO doped periodically-poled lithium niobate (ppLN) crystal is placed within a macroscopic Fabry-Pérot cavity which is doubly-resonant at 780 nm (difference-frequency or DF) and 516 nm (sum-frequency or SF) wavelengths. The 10 mm long crystal is positioned approximately at the center of the 82 mm long cavity (SI A) (b) When pumped at 1529 nm, the crystal couples the two cavities modes ($\hat{a}_{\text{DF}}$ and $\hat{a}_{\text{SF}}$) via sum- and difference-frequency generation processes. When the pump is near two-mode resonance, this coupling admixes the two modes yielding eigenmodes (with lowering operators $\hat{a}_{+}$ and $\hat{a}_{-}$) that are the superposition of the DF and SF modes. In the absence of loss, the minimum frequency splitting between the hybridized modes is determined by the coupling strength $g=\sqrt{\epsilon}\leavevmode\nobreak\ \nu$, where $\epsilon$ is the single-pass wavelength-conversion efficiency and $\nu$ is the cavity free-spectral range ($c/2L_{\text{cav}}$). (c) We measure the resonator transmission at DF and SF as a function of the detuning of the DF probe laser from the DF cavity resonance and observe a mode splitting in the spectrum when the crystal is driven by a pump. The frequency splitting is evidence of hybridization of the DF and SF modes. The solid lines are fits. The spectra correspond to a coupling $g=2\pi\times 5.0(2)$ MHz and cavity decay rates $\{\kappa_{\text{DF}},\kappa_{\text{SF}}\}=2\pi\times\{2.5(1),6(1)\}$ MHz, providing a cooperativity $\mathcal{C}=4g^{2}/\kappa_{\text{DF}}\kappa_{\text{SF}}=7(1)$.
@@ V Acknowledgments
This work was supported by AFOSR grant FA9550-22-1-0279, ARO grant W911NF-23-1-0053, AFOSR MURI grant FA9550-19-1-0399, and NSF QLCI-HQAN grant 2016136. We thank Tingran Wang and Zeyang Li for assistance with instrumentation, Ash Kumar for discussion of theory, and Zeyang Li and Adam Shaw for feedback on the manuscript.
@@ X Additional Information
Each horizontal slice on the avoided-crossing diagram in the main text figure 2 is the output of the cavity when it is probed with 780 nm light. At the two-mode resonance, we measure two peaks with equal transmission. Away from resonance, we measure more 780 nm transmission corresponding to the mode that has more contribution from the uncoupled DF eigenmode. The near-Lorentzian measurement of the off-resonant peaks is truncated in transmission due to the vertical resolution of the oscilloscope.
@@
Supplemental Information Fig S3: Estimated improvements (a) Calculation for the single-pass efficiency per mW of pump power for the pump and input beam sizes for the current cavity design (with 780 nm cavity waist $\sim$110 $\mu$m at DF), intermediate waist ($\sim$50 $\mu$m) and optimal waist ($\sim$14 $\mu$m). The maximum is obtained when the parameters $\zeta$ of the cavity and the pump are equal to 2.84, where $\zeta=L_{\text{crystal}}/2z_{R}$ and $z_{R}$ is the Rayleigh length of the pump and the cavity. The calculations are done for 1 mW of pump power using the expression in SI C. (b) Expected end-to-end conversion efficiency in different scenarios. Blue: fit corresponding to data in main text Fig 3. Orange: with optimized outcouplers with the current crystal and cavity/pump waists. Green: with optimized waists and outcouplers with the current crystal. Red: with optimized waists and outcouplers and 3x lower losses at both SF and DF. Purple: with 10x lower losses at both SF and DF. (c) Finesse measurements with an intracavity undoped, unpoled 3 mm long lithium niobate crystal provided a limiting cavity finesse of 9900 calculated from cavity finesse without crystal (7400) and cavity finesse with crystal (4200).
@@
Other crystals, such as potassium titanyl phosphate (KTP), exhibit lower absorption losses than lithium niobate (LN). However, due to their lower electro-optic coefficients (around half of LN for KTP), they require a longer interaction length. After accounting for the length, ppKTP should provide a factor of $\sim 7$ improvement in loss over 5%-MgO doped LN at 780 nm, but evaluates to comparable overall losses as compared to congruent LN (Loss of ppKTP at 775 nm has been measured to be $\sim$127 ppm per cm steinlechner2013Absorption , $\sim$3435 ppm per cm for 5%-MgO doped LN schwesyg2011Optical and 400 ppm per cm for congruently grown LN leidinger2015Comparative ).